I need to solve this problem. 4(7-3x)=1/16
Do I need to put in logarithimic form first? log4(7-3x)=log-16
Please help, I'm not sure!

Answer by Alan3354(21583)

4(7-3x)=1/16
Do I need to put in logarithimic form first? log4(7-3x)=log-16
------------
Is it 4*(7-3x) = 1/16 ?